You can move around with your keyboard
plus
minus
left
right
Drop pin on the map to start a new report
space
Press space again to adjust the location
space
Back to all reports

Problem at other roadworks problem

Reported via Open311 in the Problem at other roadworks category by Customer Services at 09:52, Mon 13 September 2021

Sent to Lincolnshire County Council less than a minute later. Council ref: 1812926.

MOP reporting that temp lights not working properly. Observed on One Network to be BT lights (work refs:
BC411LCWC8T6V6J001 and
BC411LCWC8T6V6J002)

RSS feed of updates to this problem Receive email when updates are left on this problem.

Updates

  • Thank you for your report. We will investigate and make an update to this report within 14 days.

    Posted by Lincolnshire County Council at 09:52, Mon 13 September 2021

  • State changed to: No further action

    Updated by Lincolnshire County Council at 10:02, Tue 14 September 2021

This report is now closed to updates. You can make a new report in the same location.